### v1

@@ V-A Gait Kinematics and Kinetics
Fig. 11 shows the energy expenditure using the normalized VO2 max data. The findings indicate a notable increase in metabolic cost when ambulating on sand with the unpowered exoskeleton. While the baseline controller, optimized for solid ground locomotion, did reduce the metabolic burden, the exertion remained higher compared to no exoskeleton condition. Implementing the MPC controller resulted in a $3.7$% decrease in metabolic rate relative to without wearing the exoskeleton, underscoring the efficiency of the control strategy in reducing energy cost. Paired sample t-tests revealed no statistically significant differences in metabolic cost between conditions. One possible reason for this lack of significance could be differences in the physical fitness of subjects, with some participants potentially having strong endurance or conditioning, which may have influenced the results.
@@ V-B Discussions
Analysis of knee joint angles on sand (Fig. 8) revealed that wearing the exoskeleton constrained the subject’s knee joint movement to some degrees. For instance, the unpowered and MPC-controlled modes made less flexion within the certain stance phase (i.e., 2525252540404040% of gait phase), and for Groups B, C and D, the maximum knee flexion of swing phase decreased compared with natural walking without exoskeleton. However, the MPC controller allowed for a gait more closely mimicking unassisted walking compared to other conditions. The improvement in assistance performance was evident in both the angle magnitude and gait phase.
The MPC controller significantly reduced activation in several lower-limb muscles on both terrains as summarized in Table I, including ankle extensors and flexors (TA, LG, and MG). This aligns with findings in [6] regarding increased work needed on sandy terrains to maintain COM height and forward momentum. The increase in SEM muscle activation on sand for both Groups C and D highlights the challenging nature of sandy terrain, which requires additional muscle engagement for stability but not fully offset by the exoskeletons. This observation implies for refined exoskeleton design and control strategies that can better accommodate the complex walking dynamics on sand.
@@
The knee joint generates less mechanical power than the hip and ankle but plays a key role in gait stabilization, particularly on sand. By reducing the effort at the knee during the stance phase, the exoskeleton helps distribute forces evenly across the lower limbs, improving walking efficiency. Although the knee contributes minimally to propulsion, the observed reduction in muscle activation suggests the exoskeleton effectively aids in maintaining stability. Supporting knee flexion and extension reduces the need for compensatory strategies that would otherwise heighten ankle stabilization demands. This, in turn, explains the observed decrease in TA activity during the swing phase, highlighting how knee assistance indirectly benefits the entire lower limb and contributes to improved gait efficiency. The results demonstrate that MPC consistently outperforms other conditions in reducing muscle activation and improving stability on sand, supporting the main claim that MPC offers more effective assistance than standard methods.
@@
Walking on sand requires 1.61.61.61.62.52.52.52.5 times more mechanical work and metabolic cost than on solid ground [22]. Unpowered exoskeletons exacerbated this, increasing muscle activation and metabolic cost (see Figs. 10 and 11). It was reported that knee exoskeleton did not provide a benefit in the cases where small positive mechanical work was required at the knee joint such as walking on flat surfaces [23] because of the additional mass of the exoskeleton itself. The unpowered mode and baseline controller increased metabolic cost by 9.4% and 5.4% respectively, compared to unassisted walking, aligning with previous findings of 8-15% increases on uneven terrain [1]. This increase is correlated with additional effort required to stabilize the body on sand. The unpowered exoskeleton not only adds weight but also restricts natural movement, increasing the metabolic burden. Similarly, while the baseline controller offers some assistance, it still demands extra effort to maintain balance and propulsion, which offsets the potential benefits of reduced knee muscle activation. In contrast, the MPC controller decreased metabolic cost by 3.7% comparing to the normal walking on sand without exoskeleton.
@@
A discrepancy was observed between muscle activation reduction and metabolic cost. For instance, while the baseline controller reduced activation in the RF, VM, BF, and TA muscles on sand, the overall metabolic cost increased. This may be due to the extra effort required to carry the exoskeleton and stabilize other joints, which offset the benefits of knee assistance. This underscores the need for further investigation into the relationship between muscle synergy and metabolic cost, as well as the optimization of assistive device design and control strategies, including assistance level ($\alpha$) and MPC parameters. Additionally, the current controller does not explicitly account for speed variations, and complex foot-sand interactions contribute to GRF fluctuations. We plan to explore the interplay between spatial and temporal features to further develop adaptive control strategies for various walking conditions, including different terrains and speeds, ultimately enhancing the system’s robustness and real-world applicability.
VI Conclusions
@@
This paper presented an integrated knee exoskeleton control for augmenting human locomotion on solid ground and sand. The machine learning-based GRF estimation enabled MPC for stiffness-based efficiently walking gait control on sand. The experimental findings validated the effectiveness of the proposed MPC controller in improving walking gait dynamics in both muscle activation reductions and less metabolic cost. These insights underscore the complexity of human-exoskeleton interaction on granular terrains and highlight the need for a holistic understanding of human locomotion biomechanics and energetic trade-offs. Future work includes optimizing the exoskeleton design and enhancing the control strategy to provide comprehensive assistance.

### v2

@@ 4.2 Muscle Activation and Metabolic Cost
Fig. 13 shows the energy expenditure using the normalized VO2 max data. The findings indicate a notable increase in metabolic cost when ambulating on sand with the unpowered exoskeleton. While the baseline controller, optimized for solid ground locomotion, did reduce the metabolic burden, the exertion remained higher compared to no exoskeleton condition. Implementing the MPC controller resulted in a $3.7$% decrease in metabolic rate relative to without wearing the exoskeleton, underscoring the efficiency of the control strategy in reducing energy cost. Paired sample t-tests revealed no statistically significant differences in metabolic cost between conditions.
@@ 5 Discussions
Walking on sand introduces significant biomechanical challenges, as reflected in the altered joint angle and moment profiles. Analysis of knee joint angles on sand (Fig. 9) revealed that wearing the exoskeleton constrained the subject’s knee joint movement to some degrees. For instance, the unpowered and MPC-controlled modes made less flexion within the certain stance phase (i.e., 2525252540404040% of gait phase), and for Groups B, C and D, the maximum knee flexion of swing phase decreased compared with natural walking without exoskeleton. However, the MPC allowed for a gait more closely mimicking unassisted walking compared to other conditions. The improvement in assistance performance was evident in both the angle magnitude and gait phase. The observed differences in ankle plantarflexion and hip flexion angles further emphasize the role of multi-joint coordination in navigating granular terrains. Similarly, joint moment analysis revealed a redistribution of mechanical effort across the lower limbs, especially on sand. The knee joint exhibited significantly increased extension moments throughout the stance phase under the MPC controller, highlighting its compensatory role in maintaining stability and propulsion. This redistribution aligns with the control intent of reducing reliance on hip and ankle propulsion to compensate for the COM sinkage when walking on sand by augmenting knee joint contribution.
@@
The knee joint generates less mechanical power than the hip and ankle but plays a key role in gait stabilization, particularly on sand. Fig. 9 shows a flatter knee flexion profile during the stance phase on sand, indicating the knee’s role in maintaining stability. By reducing the effort at the knee during the stance phase, the exoskeleton helps distribute forces evenly across the lower limbs, improving walking efficiency. Although the knee contributes minimally to propulsion, the observed reduction in muscle activation suggests the exoskeleton effectively aids in maintaining stability. Supporting knee flexion and extension reduces the need for compensatory strategies that would otherwise heighten ankle stabilization demands, as evidenced by the significant reduction in TA, LG, and MG muscle activations in Group D, shown in Fig. 11 and summarized in Table 1. This, in turn, explains the observed decrease in TA muscle activity during the swing phase, highlighting how knee assistance indirectly benefits the entire lower limb and contributes to improved gait efficiency. The results demonstrate that the MPC consistently outperforms other conditions in reducing muscle activation and improving stability on sand, as evidenced by the notable reductions in muscle activations in Table 1, supporting the main claim that the MPC design offers more effective assistance than standard methods.
The observed redistribution of joint moments highlights a trade-off between reducing muscle activation and managing energy expenditure. By redistributing mechanical efforts among the lower limbs, the exoskeleton reduces localized muscle fatigue. Specifically, the MPC controller significantly reduced activation in several lower-limb muscles on both terrains as summarized in Table 1, including ankle extensors and flexors (TA, LG, and MG). This aligns with findings in [10] regarding increased work needed on sandy terrains to maintain COM height and forward momentum. The increase in SEM muscle activation on sand for both Groups C and D highlights the challenging nature of sandy terrain, which requires additional muscle engagement for stability but not fully offset by the exoskeletons. In contrast, the unpowered exoskeleton (Group B) increased muscle activation due to added weight and lack of active control, reinforcing the importance of adaptive torque assistance. These observations imply that refined exoskeleton design and control strategies can accommodate the complex walking dynamics on sand.
@@
Walking on sand requires 1.61.61.61.62.52.52.52.5 times more mechanical work and metabolic cost than on solid ground [9]. Unpowered exoskeletons exacerbated high energy cost, increasing muscle activation and metabolic cost (see Figs. 12 and 13). It was reported that knee exoskeleton did not provide a benefit in the cases where small positive mechanical work was required at the knee joint such as walking on flat surfaces because of the additional mass of the exoskeleton itself [31]. The unpowered mode and baseline controller increased metabolic cost by 9.4% and 5.4%, respectively, compared to unassisted walking, aligning with previous findings of 8-15% increases on uneven terrain [1]. This increase is correlated with additional effort required to stabilize the body on sand. The unpowered exoskeleton not only adds weight but also restricts natural movement, increasing the metabolic burden. Similarly, while the baseline controller offers some assistance, it still demands extra effort to maintain balance and propulsion, which offsets the potential benefits of reduced knee muscle activation. In contrast, the MPC controller decreased metabolic cost by 3.7% comparing to the normal walking on sand without exoskeleton. Though, no statistical significance was found on the metabolic cost. One possible reason for this lack of significance could be differences in the physical fitness of subjects, with some participants potentially having strong endurance or conditioning, which may need further investigation.
@@
The above findings have significant implications for rehabilitation and assistive device design. The ability of the proposed MPC to enhance gait stability and reduce muscle activation on sand suggests potential applications for individuals with lower-limb impairments, particularly in environments with variable terrain compliance. By redistributing mechanical effort and stabilizing joint movements, the exoskeleton can improve user’s mobility and reduce muscle fatigue. On the other hand, a discrepancy was observed between muscle activation reduction and metabolic cost. For instance, the unpowered exoskeleton (Group B) significantly increased metabolic cost by 9.4%. While the baseline controller (Group C) reduced activation in the RF, VM, BF, and TA muscles on sand, the overall metabolic cost increased. These findings may be due to the extra effort required to carry the exoskeleton and stabilize other joints, which offset the benefits of knee assistance. This underscores the need for further investigation into the relationship between muscle synergy and metabolic cost, as well as the optimization of assistive device design and control strategies, including assistance level ($\alpha$) and MPC parameters. Incorporating multi-joint coordination into the control framework could further enhance stability and reduce metabolic cost by distributing effort more effectively across the lower limbs. Additionally, the current controller does not explicitly account for speed variations, and complex foot-sand interactions contribute to GRF fluctuations. The experiment setup and design are further needed for accurate GRF measurements. We plan to explore the interplay between spatial and temporal features to develop adaptive control strategies for various walking conditions, including different terrains and speeds, ultimately enhancing the system’s robustness and real-world applicability.
6 Conclusions
@@
This paper presented an integrated knee exoskeleton control for augmenting human locomotion on solid ground and sand. The machine learning-based GRF estimation enabled MPC for stiffness-based efficiently walking gait control on sand. The experimental findings validated the effectiveness of the proposed MPC controller in improving walking gait dynamics in both muscle activation reductions and less metabolic cost. The results also demonstrated that knee exoskeletons can help redistribute the kinematic and kinetic changes among the lower-limb joints. These insights underscore the complexity of human-exoskeleton interaction on granular terrains and highlight the need for a holistic understanding of human locomotion biomechanics and energetic trade-offs. Future work includes optimizing the exoskeleton design and enhancing the control strategy to provide comprehensive assistance.